OpenStudy (anonymous):

line l has equation y=-1/2x+4, and point P has coordinates (3,5). a. Find the equation of line m that passes through P and is perpendicular to l

OpenStudy (compassionate):

\[ \begin{array}l\color{red}{\text{W}}\color{red}{\text{E}}\color{red}{\text{L}}\color{red}{\text{C}}\color{red}{\text{O}}\color{red}{\text{M}}\color{red}{\text{E}}\color{red}{\text{ }}\color{red}{\text{T}}\color{red}{\text{O}}\color{red}{\text{ }}\color{red}{\text{O}}\color{red}{\text{P}}\color{red}{\text{E}}\color{red}{\text{N}}\color{red}{\text{S}}\color{red}{\text{T}}\color{red}{\text{U}}\color{red}{\text{D}}\color{red}{\text{Y}}\color{red}{\text{}}\end{array} \] Alright, so the equation can be given by plugging in the (3, 5) coordinates to the slope-intercept formula y = mx + b y = 3x + 5

OpenStudy (rajee_sam):

that is not correct Compassionate

OpenStudy (compassionate):

Oh - am I wrong? I must not have understood what he was asking.

OpenStudy (rajee_sam):

LIne I is\[y = -\frac{ 1 }{ 2} \times x + 4\] Slope is -1/2 = \[m _{1}\] A line that is perpendicular to I will have a slope \[-\frac{ 1 }{ m _{1} }\] So slope of line m is 2

OpenStudy (rajee_sam):

Now the equation of the line in Slope intercept form can be written as \[y = 2x + b\] now to find b you are given a point which says when x = 3 ; y = 5. Plug these values in the equation. \[5 = 2(3) + b\] b = -1 Now the equation of the line m which is perpendicular to Line I is \[y = 2x - 1\]